OpenStudy (anonymous):

5/2 x-1=x+ 1/4 .. 5/2 and 1/4 are fractions!

OpenStudy (anonymous):

are you asking \[ \frac{5}{2} x - 1 = x + \frac{1}{4}\] ? if so, the solution is, \[ -1 - \frac{1}{4} = x-\frac{5}{2}x \] \[-\frac{5}{4} = -\frac{3}{2}x\] \[ x = \frac{5}{6}\]

OpenStudy (anonymous):

yes! thank you